OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through points S(–1, –4) and T(3, 4). a. y = −2x − 2 c. y = 2x + 2 b. y = −2x + 2 d. y = 2x − 2

OpenStudy (nubeer):

first you have to find the gradient.. do u know how to find it?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

y=mx+c here m =4-(-4)/3-(-1) =8/4=2 y-4/x-3=2 y-4=2x-6 y=2x-2

OpenStudy (anonymous):

Thank You !